Skip to main content

AAVSOnet news 2013-09-10

1 post / 0 new
HQA's picture
AAVSOnet news 2013-09-10

The southwestern monsoon continues, so all telescopes in AZ/NM are off-line.  We expect the monsoon to break sometime in the next week or so, and operations to resume.

W30's computer was barely limping along, so Bill Goff and I decided to replace the control computer.  A new computer has been purchased (thanks, those who donated towards this project!) and is being configured by Bill.  There was also a rolloff roof problem, so no new data has come from W30 since August 26 (UT).

BSM_South is fully functional, after a computer replacement.  However, Peter Nelson is on holiday, and new data won't be taken for another couple of weeks.

OC61 is operational and taking data.  The weather is forcasted to be poor for the next few days.

BSM_Hamren (California) is nearing operation.  We've taken a couple of nights' worth of data, and will be doing final equipment tests and software configuration in a couple of weeks when Bob Stine returns to his Ranchito.

BSM_Berry (HQ) is nearing operation.  We've taken several nights of data on V339 Del.  I need to further clean the sensor and do a new polar alignment, now that Gary Walker's enclosure is in place.  I expect it to be taking survey data in about a week.

All other telescopes are in various stages of progress with no timelines available.


AAVSO 49 Bay State Rd. Cambridge, MA 02138 617-354-0484